OpenStudy (anonymous):

how do you solve this problem with either substitution or elimination? y=x+2 y=2x

OpenStudy (anonymous):

you can equal them out, you can eliminate one side or solve for x either way

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

so 2x=x+2 or y-2=x so y=2(y-2)

OpenStudy (anonymous):

oh so i make it y=2x and y=2x?

OpenStudy (anonymous):

hold on i'll write it out and attach

OpenStudy (dumbcow):

since they are both solved for y you can also just set the equations equal to each other 2x = x+2

OpenStudy (anonymous):

indeed you can solve this numerous of ways

OpenStudy (dumbcow):

yep :) always try to use the simplest way

jimthompson5910 (jim_thompson5910):

Using substitution y=x+2 2x=x+2 ... Plug in y = 2x 2x-x=2 x=2 ------------- y = 2x y = 2(2) y = 4 So the solutions are x =2 and y = 4 giving us the ordered pair (2,4)

OpenStudy (anonymous):

easiest would to just substitute the y=2x into the first equation

OpenStudy (anonymous):

thanks you guys

OpenStudy (anonymous):

right side is elimination method and your welcom

OpenStudy (anonymous):

left is substitution for x instead of y just for kicks and what not... you'll have to do it someday if you haven't yet

OpenStudy (anonymous):

k got it